OpenStudy (anonymous):

Joe packs his books in crafes that are 1.5 ft long, 10 in. high,and 14 in wide. How many cubic in. do each crate contain?

OpenStudy (anonymous):

@Xetion

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

w*l*h

OpenStudy (anonymous):

gracias:)

OpenStudy (anonymous):

1.5 ft*10 in*14 in. oh and convert all to ft or all to in

OpenStudy (anonymous):

:/ the chats dead:/

OpenStudy (anonymous):

yes,sir:)

OpenStudy (anonymous):

all to inches 1.5 ft*10 in*14 in. ft=12*in so 1.5 (12*in)*10 in*14 in=2520.0*inch^3

OpenStudy (anonymous):

1.5 ft as in is 18 in

OpenStudy (anonymous):

all to feet 1.5 ft*10 in*14 in. in=ft/12 1.5 ft*10 (ft/12)*14 (ft/12)=1.458333333*ft^3

OpenStudy (anonymous):

can i do 2520 in?

OpenStudy (anonymous):

^3?

OpenStudy (anonymous):

as my answer?

OpenStudy (anonymous):

it has to be 2520 in ^3

OpenStudy (anonymous):

yes,mk and tnx:)

OpenStudy (anonymous):

ooh i missed a part of your q. It has to be in inch so it 2520 in^3 :D

OpenStudy (anonymous):

ok,thank you,kind sir:)

OpenStudy (anonymous):

May I close the question?

OpenStudy (anonymous):

why you ask me? :P it is your q

OpenStudy (anonymous):

lol.yea.sorr- nvm lol

OpenStudy (anonymous):

I will now post another one real quick:)